OpenStudy (anonymous):

Solve by using the perfect squares method. x^2 - 4x + 4 = 0

OpenStudy (anonymous):

use perfect squares : 4x/2=2 \[2^{2}=4\] x^2 - 4x=-4 \[(x-2)^{2}=-4+4\] \[(x-2)^{2}=0\] x=2

OpenStudy (anonymous):

oh okay thank you I see what you did there and I get it now